OpenStudy (anonymous):

can someone please explain to me how to solve this. Write in slope-intercept form. m=4; (3,5)

OpenStudy (anonymous):

Algebra or Geo ?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

I can't understund what it wants. :/

OpenStudy (anonymous):

the slope is 4 and the x intercept is 3 and the y intercept is 5. put that into y=mx+b form

OpenStudy (anonymous):

y=ax+b =>5=a3+b => 5=(-4)*3+b => b=17, so,a=-4 and b=17. y=-4x+17. I think that this is the answer.